OpenStudy (anonymous):

Use the FOIL method to find the product of the binomials. (x + 7)(3x - 2) A. 3x2 + 21x - 14 B. 3x2 + 19x - 14 C. 3x2 - 2x - 14 D. 3x2 - 23x - 14

OpenStudy (anonymous):

first multiply each term ... \[\Large (x+7)(3x-2)=[x\cdot 3x]-2\cdot x+7\cdot 3x-7\cdot 2 \] now multiply\simplify them...

OpenStudy (anonymous):

i got B I think its wrong thou ..

OpenStudy (anonymous):

\[\LARGE 3x^2-2x+21x-14=3x^2+19x-14\] that's correct... why do you think it's wrong? :)